# Active robustness against the detuning-error for Rydberg quantum gates

### v1

@@ VI Discussion and Summary
We present a scheme of high-fidelity adiabatic $C_{Z}$ gates natively having active robustness to one type of errors, i.e., the Rydberg two-photon detuning error. We mainly focus on optimizing the laser pulses to be insensitive towards the deviation of two-photon detuning, which may originate from versatile experimental imperfections, such as atomic temperature effect or laser amplitude noise. Based on the use of a cleverly-modified cost function that targets at minimizing the fluctuation of infidelity(average infidelity) in a certain error range, we observe the gate error due to Doppler shifts and laser amplitude deviations can be greatly suppressed, although at the expense of a slightly lower ideal fidelity. Additionally, the optimized laser pulse can be more effective by accounting for the type of error distributions. One example is when the DER-i pulse contains Gaussian-weighted coefficients in defining the average fidelity function ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$ it can outperform other two pulses (DER and TO) with an explicitly higher gate fidelity for larger atomic temperatures (Fig.3b). This work significantly relaxes the atomic temperature constraint in a Rydberg experiment by reducing the dephasing effect from atomic velocity and position fluctuations [77]. The new DER and DER-i pulses have more advantages when the experimental imperfections are large, e.g. a high temperature $\sim$ mK, which enables the advanced atomic cooling or optical pumping techniques unnecessary for future experiments.
While we still admit that, our new pulses(DER and DER-i) do not explicitly have enhanced and active robustness to other noises as compared to the TO pulse, such as the intermediate detuning error, inhomogeneity in Rabi frequencies, laser phase noise and fluctuated interatomic interactions (see Appendix C). Especially the laser phase noise manifests as a time-dependent parameter fluctuation which contributes to the gate error at a comparable level as the laser amplitude noise [46, 64]. Finally we believe that, our results provide a more prospective method for achieving high-fidelity Rydberg gates, combining with the promising merits of active robustness to any imperfection associated with the two-photon detuning deviation. An actively noise-robust pulse will inspire further experimental improvements focusing on the suppression of one type of errors for some important tasks in the field of large-scale quantum computation [78, 79, 80, 81].
@@ Appendix A Effective dynamics for $C_{Z}$ gates
When the input state is $|01\rangle$(and $|10\rangle$) with only one excited atom, the dynamics is given by the coupling of single atom, rewritten as
@@
with the two-photon detuning $\delta^{\prime}=\delta_{0}+\delta_{ac}=\delta_{0}+\frac{\Omega_{c}^{2}-\Omega_{p}^{2}}{4\Delta_{0}}$, forming a two-level system with Rabi frequency $\tilde{\Omega}$ and detuning $\delta^{\prime}$. Whereas it becomes a little more complicated for the $|11\rangle$ state. There are six symmetric states $\{|11\rangle,\frac{|1p\rangle+|p1\rangle}{\sqrt{2}},|pp\rangle,\frac{|1r\rangle+|r1\rangle}{\sqrt{2}},\frac{|pr\rangle+|rp\rangle}{\sqrt{2}},|rr\rangle\}$ related to this problem. After adiabatic elimination of $|p\rangle$ the initial Hamiltonian for the $|11\rangle$ state is given by
@@
After eliminating $|rr\rangle$ within the Rydberg blockade regime, we arrive at an effective coupling between $|11\rangle$ and $|\chi\rangle$ as a similar two-level system
@@
Figure 5: The population dynamics and the phase evolution of states (a) $|01\rangle$ and (b) $|11\rangle$. The population returns back with the accumulated phase $\phi_{10}\approx 1.00000\pi$ and $\phi_{11}\approx 0.99902\pi$. For specific choice of the TO pulse $2\phi_{01}-\phi_{11}\approx 1.00098\pi$. Results based on the effective two-level systems(circles) have a perfect agreement with the original dynamics(lines).
Taking the specific TO pulse as an example, we numerically confirm the implementation of a two-qubit $C_{Z}$ gate by plotting the dynamical evolution of each basis state $\{|01\rangle,|11\rangle\}$ combined with their accumulated phases, see Fig.5(a-b). Via the use of complete double STIRAP pulse sequences, both $|01\rangle$ and $|11\rangle$ return to their initial positions on the Bloch sphere at $t=T_{g}$ forming an enclosed area [83]. For comparison the effective two-level dynamics is also calculated in the figure. Note that $|00\rangle$ is uncoupled and therefore, accumulates no phase. State $|01\rangle$ and $|10\rangle$ are equivalent by symmetry so $\phi_{01}=\phi_{10}$. For the case of TO pulse we clearly see the population returning back to the desired state $|01\rangle$ and $|11\rangle$ after performing the double STIRAP pulse sequences in which the total phase accumulation $\phi_{00}-\phi_{01}-\phi_{10}+\phi_{11}\approx-1.00098\pi$ realizes the $C_{Z}$ gate.
Appendix B Cost function for the DER and DER-i pulses
We now turn to study the importance of a specifically-designed cost function. To find the robust pulse, we minimize the single cost function $J_{der}$ using the Genetic optimization algorithm. The minimization of the first term $1-\mathcal{F}(\epsilon_{\delta}=0)$ is obtained when the ideal fidelity in the absence of any decay, is maximized. While the second terms $\mathcal{F}_{\max}(\epsilon_{\delta})-\mathcal{F}_{\min}(\epsilon_{\delta})$ or $1-\bar{\mathcal{F}}(\epsilon_{\delta})$ are contributed by minimizing the fluctuation of exact fidelity numbers in a certain error range $[-\epsilon_{0},\epsilon_{0}]$. In the maintext, for a set of optimal pulse parameters $(t_{1},t_{2},\omega)$, the single cost function $J_{der}$ or $J_{der-i}$ can be minimized by comparing the competing effects from two terms treating as the way of single-objective optimization.
@@ Appendix B Cost function for the DER and DER-i pulses
Figure 6: Results from the multi-objective optimization for (a) DER and (b) DER-i cases. 20 Pareto points are given standing for a good trade-off between two objective functions. Tables on the right illustrate five representative pulse parameters $(t_{1},t_{2},\omega)$(in units of $\mu$s), corresponding to the circled points in the figure.
To show the competing effect of two terms, we introduce the results based on multi-objective optimization [84], where the cost functions are composed by
@@
for the DER-i pulse. By utilizing the method of multi-objective function in the numerical Genetic algorithm, we find that, there is a very obvious competition between the two objective functions. This finding further confirms if we want to improve the pulse robustness, the fidelity at $\epsilon_{\delta}=0$(ideal fidelity) needs to be sacrificed.
In the multi-objective optimization, we introduce the Pareto solution to find the most adaptive pulse parameters. By using $J_{der1}(J_{der1-i})$ and $J_{der2}(J_{der2-i})$ as two individual objective functions, we obtain a set of Pareto points [85] as displayed in Fig.6(a) and (b) which explicitly shows an inverse relationship between them. As expected, the suppression of fidelity sensitivity to experimental imperfections requires a larger sacrifice in the ideal fidelity (for $J_{der2-i}=4.3\times 10^{-3}$ the ideal fidelity lowers to $0.9958$). Here each point presents a set of reasonable pulse parameters $(t_{1},t_{2},\omega)$ as exemplified in the Table of Fig.6. Due to the competing effect the optimal pulse parameters used for implementing, the gate should satisfy a trade-off between two aspects: ideal fidelity and robustness.
@@ Appendix C Passive influences by other noises
In order to characterize the gate performance in the presence of all experimental imperfections, in this section, we study other errors owing to the detuning deviation from the intermediate state $|p\rangle$, inhomogeneity in Rabi frequency, laser phase noise, as well as the Rydberg interaction fluctuations. We address that our robust gate is passively influenced by those errors so the DER(or DER-i) pulse does not have explicit robustness to them as compared with the TO pulse.
@@
Figure 7: The gate infidelity of the TO, DER, DER-i pulses caused by other error sources including (a) The detuning error due to intermediate-state detuning deviation $\epsilon_{\Delta}$(here $\Delta_{0}/2\pi=2.0$ GHz); (b) Inhomogeneous laser Rabi frequency coming from atom position fluctuations associated with the variation of temperature $T$; (c) Average dephasing effect denoted by rate $\gamma_{z}$ due to the time-dependent laser phase noise; (d) Random variations $\Delta B$ in the non-fluctuated interatomic interaction $B$. Each point in all figures shows an average fidelity over 500 independent realizations.
@@
Intermediate detuning error - To assess the performance of gate in a fluctuated intermediate-state detuning $\Delta_{0}\to\Delta_{0}+\epsilon_{\Delta}$, we now study the infidelity with a randomly distributed error $\epsilon_{\Delta}$. In a realistic experiment, this intermediate detuning error is inevitable due to the laser frequency fluctuation or atomic temperature effect [86]. However since a large intermediate detuning $|\Delta_{0}|\gg\Omega_{p,c}$ is set for a two-photon process; therefore, $\epsilon_{\Delta}\ll|\Delta_{0}|$ which makes the detuning error negligible. We numerically verify the negligible effect of $\epsilon_{\Delta}$ on the gate infidelity $\mathcal{F}(0)-\mathcal{F}(\epsilon_{\Delta})$ for all three pulses as shown in Fig.7(a). Note that the fluctuated detuning $\Delta_{0}+\epsilon_{\Delta}$ is also reversed to $-(\Delta_{0}+\epsilon_{\Delta})$ halfway for cancelling the dynamical phase. It is clear that the infidelity caused by different $\epsilon_{\Delta}$ values stays below $10^{-6}$. This is expected because the large intermediate detuning $\Delta_{0}$ can efficiently suppress the intermediate population during the gate duration leading to a negligible error [87]. In addition we observe the TO pulse shows a relatively better robustness $\sim 10^{-8}$ to the $\epsilon_{\Delta}$ because of the presence of a perfect dark state $|d\rangle$(see Sec. III.A). While the DER or DER-i pulse is more robust for the deviation of $\epsilon_{\delta}$, not necessarily the most robust for the $\epsilon_{\Delta}$. Finally since $\Delta_{0}$ is not an optimal value, we also observe the infidelity oscillations with respect to the change of $\epsilon_{\Delta}$.
@@
Inhomogeneous Rabi frequency - In general the atom position fluctuations are inevitable in experiments which directly lead to a position-dependent laser Rabi frequency. Therefore the realistic Rabi frequency perceived by the atoms will deviate from its ideal value. According to Ref.[88], we replace $\Omega_{p}(t)$ and $\Omega_{c}(t)$ by position-dependent functions $\Omega_{p}(\mathbb{r},t)$ and $\Omega_{c}(\mathbb{r},t)$, which are expressed as
@@
where $\Omega_{p,c}(0,t)=\Omega_{p,c}(t)$ is the originally optimized waveform and $\mathbb{r}=(x,y,z)$ presents the atom position at any time $t$ that satisfies a Gaussian distribution
@@
When the laser waists are $\omega_{x,p}=\omega_{y,p}=7.8$ $\mu$m, $\omega_{x,c}=\omega_{y,c}=8.3$ $\mu$m, arising the Rayleigh lengths $L_{x(y),p}=455.08$ $\mu$m, $L_{x(y),c}=213.65$ $\mu$m, respectively for the 420 nm and 1013 nm lasers. The standard deviation $\sigma_{x,y,z}$ in $f(\mathbb{r})$ can be estimated by $\sigma_{x,y,z}=\sqrt{k_{B}T/m\omega_{x,y,z}^{2}}$ with trapping frequencies $\omega_{x,y,z}/2\pi=(147,117,35)$ kHz chosen from [89], which arises the maximal uncertainty in the atom position $(\sigma_{x},\sigma_{y},\sigma_{z})=(0.47,0.60,1.99)$ $\mu$m at $T=2$ mK. That means the realistic deviation of Rabi frequency felt by atoms is still very small owing to $\sigma_{x}\ll\omega_{x,p(c)}$, $\sigma_{y}\ll\omega_{y,p(c)}$ even at a high temperature.
Our results are summarized in Fig.7(b) where the infidelity caused by inhomogeneous Rabi frequency varies with the atomic temperature $T$. We observe that, although the waist of a Gaussian laser can be adjustable experimentally by using different incident angles, for a reasonable temperature, atoms in the optical traps can still be uniformly illuminated by both Rydberg excitation lasers [90]. Because the maximal deviation of atom position is smaller than the beam waist by orders of magnitude. From Fig.7(b) the robust pulses do not have explicit advantages against this inhomogeneity error and the DER and TO pulses reveal a similar trend of insensitivity. Here we note that the DER-i pulse shows an anomalous reduction in the gate infidelity achieving $|\mathcal{F}(0)-\mathcal{F}(T)|\approx 1.31\times 10^{-5}$ for $T=2$ mK. This is because the originally optimized pulse $\Omega_{p,c}(0,t)$ is explicitly robust to the two-detuning error, but not the best to other errors such as the error induced by inhomogeneous Rabi frequency. Slightly shifting Rabi frequency via atom position can instead improve the gate insensitivity to the variation of temperature.
Laser phase noise - The Rydberg excitation lasers, not only have amplitude fluctuation as discussed in Sec.VB, but also contribute significant phase noise. Due to different frequencies involved in excitation lasers, the realistic Rabi frequency can be described by
@@
where $\phi_{p,c}(t)$ presents a random phase typically featuring time-dependent fluctuations [64].
@@
represent dephasings of the control and target atoms with individual rates $\gamma_{1},\gamma_{2}$ associated with the optical transitions of $|p\rangle\to|1\rangle$ and $|r\rangle\to|p\rangle$, which are randomly obtained from the range of $[0,\gamma_{z}]$. Figure 7(c) shows the gate infidelity caused by the increase of $\gamma_{z}$. Due the impact of laser phase noise we see all three cases reveal a similar trend with an exponential enhancement in the gate infidelity $\mathcal{F}(0)-\mathcal{F}(\gamma_{z})$. When $\gamma_{z}/2\pi=40$ kHz we find the average infidelity even reaches $\sim 10^{-2}$ after running 500 simulations.
Fluctuation of Rydberg interaction - Our scheme relies on the Rydberg-blockade mechanism that are robust to the fluctuation of interatomic interaction [92]. Under finite temperature, the atom position in a trap will lead to a fluctuation $\Delta B$ in interaction strength deviating from its expected value $B$. While benefiting from a strong long-range interaction to suppress the excitation of $|rr\rangle$, we find the impact of interaction fluctuation is very small. As displayed in Fig.7(d), we calculate the infidelity of the three pulses by varying the ratio $\Delta B/B$ in the range of $[0,0.1]$. The real fluctuation is a random number obtained from $[-\Delta B,\Delta B]$ for each realization. If $\Delta B/B=0.1$($B/2\pi=2.0$ GHz), the relative position deviation between two atoms is about $\delta r=r_{0}^{7}\Delta B/(6C_{6})\approx 45.83$ nm [93], estimated by the position-dependent interaction strength $B=C_{6}/r_{0}^{6}$ with $C_{6}=862.69$ $\rm{GHz}\cdot\mu m^{6}$($70S_{1/2}$) and $r_{0}=2.75$ $\mu$m [94]. Here, the infidelity of all pulses keeps at a very low level $10^{-12}\sim 10^{10}$ offering a negligible impact on the gate operations.

### v2

@@ VI Discussion and Summary
We present a scheme of high-fidelity adiabatic $C_{Z}$ gates natively having active robustness to one type of errors, i.e., the Rydberg two-photon detuning error. We mainly focus on optimizing the laser pulses to be insensitive towards the deviation of two-photon detuning, which may originate from versatile experimental imperfections, such as atomic temperature effect or laser amplitude deviation. Based on the use of a cleverly-modified cost function that targets at minimizing the fluctuation of infidelity(average infidelity) in a certain error range, we observe that, the gate error due to Doppler shifts and laser amplitude deviations can be reduced, although at the expense of a slightly lower ideal fidelity at small detunings. Additionally, the optimized laser pulse can be more effective by accounting for the type of error distributions. One example is when the DER-i pulse contains Gaussian-weighted coefficients in defining the average fidelity function, it can outperform other two pulses (DER and TO) with an explicitly higher gate fidelity for lower atomic temperatures. In addition, the newly-proposed DER and DER-i pulses under a fine optimization have higher gate fidelities for any temperature, which significantly relaxes the atomic temperature constraint in a Rydberg experiment by reducing the dephasing effect from atomic velocity and position fluctuations [79].
While we still admit that, the robust pulses (DER and DER-i) do not explicitly have enhanced and active robustness to other types of noises, such as the intermediate detuning error, inhomogeneity in Rabi frequencies, laser amplitude and phase noises, and the fluctuated interatomic interactions (see Appendix C). Especially the laser phase noise manifests as a time-dependent parameter fluctuation which contributes to the gate error at a comparable level as the laser amplitude noise [47, 68]. Finally we believe that, our results provide a more prospective method for achieving high-fidelity Rydberg gates, combining with the promising merits of active robustness to any imperfection associated with the two-photon detuning deviations. An actively noise-robust pulse will inspire further experimental improvements focusing on the suppression of one type of errors for some important tasks in the field of large-scale quantum computation [80, 81, 82, 83].
@@ Appendix A Effective dynamics for $C_{Z}$ gates
Figure 5: The population dynamics associated with the effective two-level model and the corresponding phase evolution of initial states (a1) $|01\rangle$ and (a2) $|11\rangle$ with the TO pulse, where no decay is considered. Similarly, (b1-b2) present the case with the DER pulse for $\epsilon_{0}/2\pi=0.8$ MHz. Other intermediate-state population is not shown. For the TO pulse, the population returns back with the accumulated phases $\phi_{10}\approx 1.00000\pi$ and $\phi_{11}\approx 0.99999\pi$, resulting in $2\phi_{01}-\phi_{11}\approx 1.00001\pi$. Analogously, for the DER pulse, $\phi_{10}\approx 1.00000\pi$ and $\phi_{11}\approx 0.99902\pi$ arising $2\phi_{01}-\phi_{11}\approx 1.00098\pi$. Good agreement is observed between the effective two-level systems(circles) and the originally full dynamics(lines).
@@
When the input state is $|01\rangle$(equally for $|10\rangle$) with only one excited atom, the dynamics is given by the coupling of single atom, rewritten as
@@
with the two-photon detuning $\delta^{\prime}=\delta_{0}+\delta_{ac}=\delta_{0}+\frac{\Omega_{c}^{2}-\Omega_{p}^{2}}{4\Delta_{0}}$, forming a two-level system with Rabi frequency $\tilde{\Omega}$ and detuning $\delta^{\prime}$.
Whereas, it becomes a little more complicated for the $|11\rangle$ state. There are six symmetric states $\{|11\rangle,\frac{|1p\rangle+|p1\rangle}{\sqrt{2}},|pp\rangle,\frac{|1r\rangle+|r1\rangle}{\sqrt{2}},\frac{|pr\rangle+|rp\rangle}{\sqrt{2}},|rr\rangle\}$ related to this problem. After adiabatic elimination of $|p\rangle$ the initial Hamiltonian for state $|11\rangle$ can be described by
@@
After eliminating $|rr\rangle$ within the strong Rydberg blockade regime, we arrive at an effective coupling between $|11\rangle$ and $|\chi\rangle$ as a similar two-level system
@@
Figure 6: Results from multi-objective optimization for the robust DER and DER-i pulses, under different detuning deviations $\epsilon_{0}/2\pi=0.8$ MHz (right panels) and 0.3 MHz (left panels). Sufficient Pareto points are found standing for a good trade-off between two objective functions. Tables on the right illustrate four representative pulse parameters $(t_{1},t_{2},\omega_{1},\omega_{2})$(in units of $\mu$s), corresponding to the circled points in the figure. The last row (highlighted in dark green) presents the optimal parameters based on single-objective function, same as in Table I.
To verify the effective two-level dynamics, we now numerically plot the implementation of a two-qubit $C_{Z}$ gate by taking the specific TO and DER pulses as examples. In Fig.5(a1-a2), the dynamical evolutions of initial $|01\rangle$ and $|11\rangle$ states, the intermediate $|0r\rangle$ and $|\chi\rangle$ states, combined with their accumulated phases $\phi_{01},\phi_{11}$, are presented for the TO pulse. Remarkably, via the use of double STIRAP pulse sequences(see top), both $|01\rangle$ and $|11\rangle$ return to their initial positions on the Bloch sphere forming a completely enclosed area [85]. This implies that the sufficient population exchange between $|01\rangle$(or $|11\rangle$) and $|0r\rangle$(or $|\chi\rangle$) appears by obeying the full adiabatic transfer and other intermediate-state population has been strongly suppressed. The resulting infidelity achieves $1-\mathcal{F}_{to}\approx 3\times 10^{-7}$ at zero detuning (see Table I). Here, note that $|00\rangle$ is uncoupled and therefore, accumulates no phase. State $|01\rangle$ and $|10\rangle$ are equivalent by symmetry so $\phi_{01}=\phi_{10}$. Hence, we clearly see the population returning back to the desired state $|01\rangle$ and $|11\rangle$ after performing the double STIRAP pulse sequences in which the total phase accumulation is $\phi_{00}-\phi_{01}-\phi_{10}+\phi_{11}\approx-1.00001\pi$ which realizes the $C_{Z}$ gate.
For comparison, we also study the population dynamics for the DER pulse in Fig. 5(b1-b2). We remark that the DER pulse requires a longer time-spent on intermediate Rydberg states $|0r\rangle$ and $|\chi\rangle$, giving rise to a less sufficiently adiabatic transfer due to the staying of Rydberg population. In this case the DER pulse becomes inherently more robust against the detuning errors on Rydberg states, and achieves $1-\mathcal{F}_{der}\approx 2.9\times 10^{-3}$ when the detuning deviation is zero. Based on the Bloch sphere version we also clearly see the adiabatic transfer is incomplete for the $|01\rangle$ and $|11\rangle$ states, and hence confirm that the TO pulse can outperform the DER pulse by having a smaller infidelity at zero detuning.
Appendix B Balanced optimization for DER and DER-i pulses
We now turn to study the importance of a specifically-designed cost function that can balance the competing effects from ideal fidelity and gate robustness. To find such a robust pulse, we minimize the single cost function using the Genetic optimization algorithm. See Eq.(14) and (16), the minimization of the first term $1-\mathcal{F}(\epsilon_{\delta}=0)$ is obtained when the ideal fidelity in the absence of any decay, is maximized. While the second terms $\mathcal{F}_{\max}(\epsilon_{\delta})-\mathcal{F}_{\min}(\epsilon_{\delta})$ or $1-\bar{\mathcal{F}}(\epsilon_{\delta})$ are contributed by minimizing the fluctuation of exact fidelity numbers in a certain error range $[-\epsilon_{0},\epsilon_{0}]$. In the maintext, based on a set of optimal pulse parameters $(t_{1},t_{2},\omega_{1},\omega_{2})$, a single cost function $J_{der}$ or $J_{der-i}$ has been minimized by comparing the competing effects from two terms, so-called the single-objective optimization.
In order to show the competing effect of two terms which arises a clear balance between the realistic gate fidelity and the robustness to detuning deviations, here we discuss the results making use of multi-objective optimization [86], where the cost functions are replaced by
@@ Appendix B Balanced optimization for DER and DER-i pulses
for the DER-i pulse. By utilizing the multi-objective function in the numerical Genetic algorithm, we observe that, there is a very obvious competition between the two objective functions. This finding further confirms that if we want to improve the pulse robustness, the ideal fidelity at small deviations needs to be sacrificed.
In the multi-objective optimization, we introduce the Pareto solution to find the most adaptive pulse parameters. By using $J_{der1}(J_{der1-i})$ and $J_{der2}(J_{der2-i})$ as two individual objective functions, we obtain a set of Pareto points [87] as displayed in Fig.6, which explicitly shows an inverse relationship between them. As expected, the suppression of infidelity fluctuation (vertical axis) to experimental imperfections requires a larger sacrifice in the ideal fidelity (horizontal axis). E.g. for the DER0.8 pulse, if we require the ideal infidelity close to $10^{-6}$(point 1) the fidelity fluctuations have to stay at a very high level $\sim 1.2\times 10^{-2}$ indicating a poor robustness to the detuning errors. But, when we could endure a relatively high gate infidelity $1-\mathcal{F}_{der}\approx 2.8\times 10^{-3}$ (point 4), then the insensitivity can be lowered to $\sim 5.4\times 10^{-3}$.
However, depending on a fine optimization, the DER-i0.3 pulse is able to achieve a higher fidelity at $\epsilon_{\delta}=0$, while still keeping strong robustness to the detuning errors. E.g. here if $1-\mathcal{F}_{der-i}\approx 1.2\times 10^{-5}$ (point 1) the average insensitivity stays at $1.1\times 10^{-4}$. In contrast if $1-\mathcal{F}_{der-i}\approx 3.6\times 10^{-5}$ (point 4) the insensitivity to detuning errors can be as small as $8.9\times 10^{-5}$. Hence, due to the competing effect the optimal pulses used for implementing the gate should satisfy a trade-off between two aspects: ideal fidelity and robustness. However, we note that the fine optimization procedure is able to provide greatly-improved gate performance maintaining a high gate fidelity as well as actively strong robustness against the detuning errors.
@@ Appendix C Passive influences by other noises
In order to characterize the gate performance in the presence of all experimental imperfections, in this section, we study other errors owing to the detuning deviation from the intermediate state $|p\rangle$, inhomogeneity in Rabi frequency, laser phase and amplitude noises, as well as the Rydberg interaction fluctuations. We address that our robust gate is passively influenced by those errors so the DER and DER-i pulses do not have explicit robustness to them as compared to the TO pulse.
@@
Figure 7: The gate infidelity of the TO, DER and DER-i (Gaussian-weighted) pulses caused by other error sources including (a) The detuning error due to the intermediate detuning deviation $\epsilon_{\Delta}$(note that the non-fluctuated part is $\Delta_{0}/2\pi=2.0$ GHz); (b) Inhomogeneous laser Rabi frequency coming from atomic position fluctuations associated with the variation of temperature $T$; (c) Average dephasing effect denoted by rate $\gamma_{z}$ due to the time-dependent laser phase noise; (d) Laser amplitude error caused by the amplitude deviation $\epsilon_{\Omega}$. Each point in all figures is obtained by averaging over 500 independent realizations.
Intermediate detuning error - To assess the performance of gate in a fluctuated intermediate-state detuning $\Delta_{0}\to\Delta_{0}+\epsilon_{\Delta}$ (also see Eq.2), we now study the gate infidelity with a randomly distributed error $\epsilon_{\Delta}$. In a realistic experiment, this intermediate detuning error is inevitable due to the laser frequency fluctuation or atomic temperature effect [88]. However, since a large non-fluctuated intermediate detuning $|\Delta_{0}|\gg\Omega_{p,c}$ is set for a two-photon process; therefore, $\epsilon_{\Delta}\ll|\Delta_{0}|$ which could make this detuning error negligible.
@@
We numerically verify the negligible effect of $\epsilon_{\Delta}$ on the absolute gate infidelity $|\mathcal{F}(0)-\mathcal{F}(\epsilon_{\Delta})|$ for all pulses as shown in Fig.7(a). Note that the fluctuated detuning $\Delta_{0}+\epsilon_{\Delta}$ is also reversed to $-(\Delta_{0}+\epsilon_{\Delta})$ halfway for cancelling the dynamical phase. It is clear that the infidelity caused by different $\epsilon_{\Delta}$ values stays below $10^{-6}$. This is expected because the large intermediate detuning $\Delta_{0}$ can efficiently suppress the intermediate population during the gate duration leading to a negligible error [89]. In addition we observe all pulses except DER0.8 reveal comparably better robustness $<10^{-7}$ to the $\epsilon_{\Delta}$ because of the presence of dark state $|d\rangle$(see Sec. III.A). The DER0.8 suffers from a slightly larger infidelity so the intermediate detuning error turns to be the worst. While the DER and DER-i pulses have shown their improved robustness against the deviation of $\epsilon_{\delta}$, not necessarily the most robust for the $\epsilon_{\Delta}$. Finally since $\Delta_{0}$ is not an optimal value, we also observe the infidelity oscillations with respect to the change of $\epsilon_{\Delta}$.
@@
Inhomogeneous Rabi frequency - In general the position fluctuations of atoms are inevitable in experiments, which directly leads to a position-dependent laser Rabi frequency. Therefore the realistic Rabi frequency perceived by the atoms will deviate from its desired value. According to Ref.[90], we replace $\Omega_{p}(t)$ and $\Omega_{c}(t)$ by position-dependent functions $\Omega_{p}(\mathbb{r},t)$ and $\Omega_{c}(\mathbb{r},t)$, described by
@@
where $\Omega_{p,c}(0,t)=\Omega_{p,c}(t)$ is the originally optimized waveform (see Eqs. 3-4) and $\mathbb{r}=(x,y,z)$ presents the atom position at any time $t$ that satisfies a Gaussian distribution
@@
When the laser waists are $\omega_{x,p}=\omega_{y,p}=7.8$ $\mu$m, $\omega_{x,c}=\omega_{y,c}=8.3$ $\mu$m, arising the Rayleigh lengths $L_{x(y),p}=455.08$ $\mu$m, $L_{x(y),c}=213.65$ $\mu$m, respectively for the 420 nm and 1013 nm lasers. The standard deviation $\sigma_{x,y,z}$ in $f(\mathbb{r})$ can be estimated by $\sigma_{x,y,z}=\sqrt{k_{B}T/m\omega_{x,y,z}^{2}}$ with trap frequencies $\omega_{x,y,z}/2\pi=(147,117,35)$ kHz chosen from [91], which arises the maximal uncertainty in the atom position $(\sigma_{x},\sigma_{y},\sigma_{z})\approx(0.47,0.60,1.99)$ $\mu$m at $T=2$ mK. That means the realistic deviation of Rabi frequency felt by atoms is still very small owing to $\sigma_{x}\ll\omega_{x,p(c)}$, $\sigma_{y}\ll\omega_{y,p(c)}$ even at a high temperature.
Our numerical results are summarized in Fig.7(b), where the infidelity caused by inhomogeneous Rabi frequency varies with the atomic temperature $T$. We observe that, although the waist of a Gaussian laser can be adjustable experimentally by using different incident angles, for a reasonable temperature, atoms in the optical traps can still be uniformly illuminated by both Rydberg excitation lasers [92]. Because the maximal deviation of atom position is smaller than the beam waist by orders of magnitude. From Fig.7(b) the robust pulses do not have explicit advantages against this inhomogeneity error, and the finely-optimal DER, DER-i pulses reveal an exponential increase in the gate infidelity, analogously to the TO pulse. This is because the originally optimized pulse $\Omega_{p,c}(0,t)$ is explicitly robust to the two-detuning error, but not the best to other errors such as the Rabi inhomogenity. Slightly shifting Rabi frequency via atom position could instead improve the gate insensitivity to the variation of temperature, e.g. for the DER-i pulse under a wide optimization.
Noises from excitation lasers - The Rydberg excitation lasers will not only contribute significant phase noise, but also have amplitude fluctuations. The former could change the realistic Rabi frequency by adding phase terms as
@@
due to different frequencies involved in the excitation lasers. Here, $\phi_{p,c}(t)$ presents a random phase typically featuring time-dependent fluctuations [68].
@@
represent dephasings of the control and target atoms with individual rates $\gamma_{1},\gamma_{2}$ associated with the optical transitions of $|p\rangle\to|1\rangle$ and $|r\rangle\to|p\rangle$, which are randomly obtained from the range of $[0,\gamma_{z}]$ ($\gamma_{z}$ treats as the maximal dephasing rate). Figure 7(c) shows the gate infidelity caused by the increase of $\gamma_{z}$. Due the impact of laser phase noise we see all pulses (including non-robust and robust pulses) exhibit a same trend with an exponential enhancement in the gate infidelity $\mathcal{F}(0)-\mathcal{F}(\gamma_{z})$. When $\gamma_{z}/2\pi=40$ kHz we find the average infidelity even reaches $\sim 10^{-2}$ after running 500 simulations.
Furthermore, the latter can be described by an uncertainty in Rabi frequencies
To distinguish from the ac Stark shift error which is also induced by the amplitude deviation (see Sec. V.B.), here we consider the detuning error $\epsilon_{\delta}=0$ and modify the sing-qubit Hamiltonian (2) as
[…]
-- $\Delta_{0}|p\rangle_{j}\langle p|$
Assuming the amplitude deviations $\epsilon_{\Omega_{p}},\epsilon_{\Omega_{c}}$ are random within the range of $[-\epsilon_{\Omega},+\epsilon_{\Omega}]$ we calculate the infidelity $\mathcal{F}(0)-\mathcal{F}(\epsilon_{\Omega})$ as a function of the amplitude error $\epsilon_{\Omega}$ in Fig.7(d). It is clearly seen that the laser amplitude noise plays a less significant impact as compared to the phase noise (the former), maintaining smaller than $2.5\times 10^{-3}$ for large values of deviation $\epsilon_{\Omega}$ up to 0.05, which is even regardless of the pulse types.
Fluctuation of Rydberg interaction - Our scheme relies on the Rydberg blockade mechanism that are robust to the fluctuation of interatomic interaction [94]. Under finite temperature, the atom position in a trap will lead to a fluctuation $\Delta B$ in the interaction strength deviating from its expected value $B$. While benefiting from a strong long-range interaction to suppress the excitation of $|rr\rangle$, we find the impact of interaction fluctuation is very small. A rough estimation shows that, if $\Delta B/B=0.1$($B/2\pi=2.0$ GHz), the relative position deviation between two atoms is about $\delta r=r_{0}^{7}\Delta B/(6C_{6})\approx 45.83$ nm [95], estimated by the position-dependent interaction strength $B=C_{6}/r_{0}^{6}$ with $C_{6}=862.69$ $\rm{GHz}\cdot\mu m^{6}$($70S_{1/2}$) and $r_{0}=2.75$ $\mu$m [96]. Then the gate infidelity of all pulses stays below $10^{-10}$(not shown) providing a a negligible impact on the gate operations.